How many 4-digit whole numbers using the digits in the set {1,2,3,4,5,6,7,8,9,0}can be formed using the restrictions below.
For purposes of this problem, the whole number can be "padded" with zeros on the left. So a 0014 is a good number,(IE numbers can start with zeros).
Each part below is is independant of the previous part.
a. Numbers do not contain the same digit twice, (without replacement)
b. Numbers end with an even digit and can repeat digits, (with replacement)
c. Numbers begin with an odd digit and can repeat digits, (with replacement)
d. Numbers begin and end with a 1, (with replacement)
e. Numbers have exactly three digits which are 7s (with replacement)

Answer by Edwin McCravy(20054)

How many 4-digit whole numbers using the digits in the set {1,2,3,4,5,6,7,8,9,0}can be formed using the restrictions below.
For purposes of this problem, the whole number can be "padded" with zeros on the left. So a 0014 is a good number,(IE numbers can start with zeros).
Each part below is is independant of the previous part.

a. Numbers do not contain the same digit twice, (without replacement)
Examples would be 0692, 8591

Choose digit #1 10 ways, and then choose digit #2 9 ways, 
and then choose digit #3 8 ways, and then choose digit #4
7 ways.There are 10×9×8×7 = 5040 ways. 

b. Numbers end with an even digit and can repeat digits, (with replacement)
Examples would be 8736, 4994, 7770, 0000, 2010

We do not have to choose the digits in any certain order. 
So we will choose digit #4, then digit #1, then digit #2,
and finally digit #3.

Choose digit #4 5 ways (2,4,6,8, or 0), and then choose 
digit #1 10 ways, and then choose digit #2 10 ways, and 
then choose digit #3 10 ways. There are 
5×10×10×10 = 5000 ways.

c. Numbers begin with an odd digit and can repeat digits,
(with replacement)
Examples:  1887, 9742, 5555

Choose digit #1 5 ways (1,3,5,7,9), and then choose 
digit #2 10 ways, and then choose digit #3 10 ways, and 
then choose digit #4 10 ways.  There are 
5×10×10×10 = 5000 ways.

d. Numbers begin and end with a 1, (with replacement)
Examples:  1871, 1661, 1111

Choose digit #1 1 way, and then choose digit #2 10 ways, 
and then choose digit #3 10 ways, and then choose digit 
#4 1 way. There are 1×10×10×1 = 100 ways.

e. Numbers have exactly three digits which are 7s (with replacement)
Examples: 7771, 7707, 7477, 0777

There are three 7's and one non-7.  We choose the position 
for the non-7 any of 4 ways, and then the digit for the non-7
9 ways, then 1 way each for the 7's.
There are 4×9×1×1 = 36 ways. 

Edwin